OpenStudy (anonymous):

How do I do this Evaluate the expression for the given value of the variable(s). \(\frac{-(4h - 1)}{1 + h} ; h = 2 \)

OpenStudy (anonymous):

substitute

OpenStudy (anonymous):

-(4h - 1)/1+ h = -(4*2 - 1)/1+1 = -(8-1)/2 = - 7/2

OpenStudy (anonymous):

sorry , denominator is 1 + 2 = 3 , so answer is -7/3

OpenStudy (anonymous):

gettit?

OpenStudy (nirmalnema):

simply replace the value of h in the ques.....

OpenStudy (nirmalnema):

-4(2)+1 / 1+(2) -7/3

OpenStudy (nirmalnema):

@SamanthaGreenfield did u understand